OpenStudy (anonymous):

4. You deposit $200.00 into an account that earns 5% interest annually. How much money will you have after 2 years?

OpenStudy (phi):

one way is figure out 200* 1.05^2 5% is 0.05, the first year you get 200+ 200*0.05 (principal plus interest) or 200(1+0.05) or 200*1.05 the 2nd year you get 200*1.05 + (200*1.05)*0.05= 200*(1.05+1.05*0.05)= 200* 1.05(1+0.05) = 200* 1.05*1.05= 200*1.05^2